OpenStudy (anonymous):

how do you solve equations like X/4 = 3x-5

OpenStudy (anonymous):

OK lets take the this equation. How will you bring 3x to the left side?

OpenStudy (anonymous):

@kikeys0419 are you there?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }x=3x -5\] \[\frac{ 1 }{ 4 }x -3x = -5\] \[\frac{-11}{4}x=-5\] \[x=\frac{20}{11}\]

OpenStudy (anonymous):

Or, multiply both sides by 4 first to clear the fractions.